# Hierarchical Reinforcement Learning for Swarm Confrontation with High Uncertainty

### v1

@@ V-A Setting Up
Before analyzing the comparative results, we first introduce the detailed settings of simulations. Following the convention in [10, 25, 37], we generate the starting locations of pursuers and evaders in the rectangular area ($D_{1}=40m,D_{2}=20m,D_{3}=30m$). The target point of evaders is behind the pursuers. Three scenarios including ten pursuers versus ten evaders (termed as V10101010), fifteen pursuers versus fifteen evaders (termed as V15151515), and twenty pursuers versus twenty evaders (termed as V20202020) are considered in our experiment, all in the presence of four moving obstacles. In each scenario, we set up five different abilities of pursuers and evaders. The capture radius of pursuers $\rho_{c}$ is equalized from $0.6m$ to $1m$. The maximum velocities of pursuers are all set to $0.5m/s$. The maximum velocities of evaders are equalized from $0.6m/s$ to $1m/s$. The radius of the agents and obstacles are $0.2m$ and $0.5m$, respectively.
@@ V-C Comparison Analysis in Cross–Training
We further deploy the policy networks trained by the learning–based methods in $100$ different instances and compare them with the non–learning algorithms. We propose two additional evaluation metrics: decision time and confrontation win rate. Decision time refers to the total time taken by all pursuers for task allocation and path planning based on their observations. We take the average value after running different instances. The confrontation win rate refers to the ratio of the number of pursuers’ successes to the number of all instances. The experiment results of $100$ instances in different sizes are shown in Table II. The table gathers the episode returns (Re.), decision time (Ti.), and confrontation win rate (W.R.) of all methods. We can observe that there is a relationship between the win rate and episode returns, which is calculated by the reward function in this study. The higher the returns achieved by pursuers, then the greater the win rate will be. Among these three non–learning algorithms, the heuristic algorithm has a longer decision time, although its returns and win rate are higher as the problem size increases. Due to the end–to–end approach, DRL has the shortest decision time, but its returns and win rate are much less than HRL. IMVE enhances the training efficiency of the method, but there is no additional improvement in its decision–making performance in the case of policy convergence. As a result, all three metrics for HRL and HRL/IMVE are similar in different scales. The results show that our method achieves better episode returns, decision time, and confrontation win rate than the baselines, especially on large–scale instances. The reason for this is that swarm confrontation in a dynamic obstacle environment has a high degree of uncertainty, and the baselines lack the ability to handle it. Our method optimizes the dynamic mechanism between the two layers based on the probabilistic ensemble model, which quantifies the uncertainty in the scenario.
@@ V-D Ablation Study for Adaptive Interaction Frequency
Moreover, we plot the value of $H$ for the adaptive method in different sizes and for different episodes (0060606060) of cross–training in Fig. 7. The horizontal coordinate refers to the time steps in each episode, and the vertical coordinate refers to the value of $H$. In the early stage of training, the method decreases the interaction step $H$ to augment the training samples in the upper layer. Through constant training, the epistemic uncertainty of the environment model decreases, so the interaction frequency becomes progressively lower. However, since aleatoric uncertainty cannot be eliminated, the increment in $H$ stabilizes after episode $15$. In addition, during time steps 50505050250250250250, the interaction frequency will be higher than the other time steps because the pursuers will meet with evaders and obstacles, and the aleatoric uncertainty will be higher at this time. Based on the adaptive frequency method, we can overcome the negative effects of uncertainty on cross–training, and finally achieve a favorable training effect.
@@ V-E Ablation Study on ITM
Figure 9: Generalization to larger–scale instances.
@@ V-F Generalization to Larger–Size Swarms
We compare the model trained in V20 with the baselines in scenarios V25–V35 similar to Table II. The experiment results are shown in Table III, which displays the episode returns (Re.), decision time (Ti.), and confrontation win rate (W.R.) testing in different problem sizes. As can be seen, HRL outperforms other baselines in all instances. To further investigate the generalization ability of our method, we utilize the trained model to solve instances with eight obstacles in Table III. An increase in the number of obstacles leads to a decrease in path rewards, but has a very small effect on confrontation win rate. The results demonstrate that the trained policies under small–size swarms are able to deal with the uncertainty in various scale scenarios. Therefore, our method has a strong generalization ability on the swarm confrontation problem.
@@ V-G Emerged Coordinated Confrontation Behaviors
With sufficient training from scratch, the proposed HRL method can emerge effective target allocation and path planning strategies. In this section, we visualize the results of trained networks with our method on one of the V10 instances in Fig. 10, and analyze the coordinated behaviors in each stage. Pursuers and evaders are labeled with "$X$" indicating their identification numbers. The figure illustrates the process of target allocation and path planning by pursuers through policy networks at different time steps. In the figure, the upper layer assigns targets based on pursuers’ and evaders’ attributes to maximize the total reward in (25). The lower layer further plans the safe paths to chase evaders while avoiding collisions with obstacles and neighbors.
@@
Time steps 50505050100100100100: After time step $50$, the pursuers begin to encounter evaders and obstacles, introducing more environment uncertainty into decision making. Since pursuers with a large radius have a wider capture range, the upper layer will prioritize mobilizing them to chase evaders with faster escape velocities, such as evader $4$ and evader $7$.
@@
Time steps 200200200200250250250250: After the evaders have escaped the initial round–up by the pursuers, the latter will choose to turn around and chase. The upper layer will assign the pursuers priority to evaders who are closer because they have a higher probability of being captured. Evaders may be hindered by an obstacle or a neighbor ahead, resulting in a loss of velocities, and the pursuers eventually catch up with them.
Time steps 250250250250300300300300: Even though evader $9$ is approaching the target area, the others have been captured by pursuers. According to the rules in Section III, pursuers capture more than half of the evaders, which win the confrontation. Therefore, our trained models could deliver reasonably good solutions in the swarm confrontation.
@@ VI Conclusion
This paper has presented a guaranteed stable HRL method for hybrid decision spaces and high uncertainty in swarm confrontation. It has designed two–layer DRL networks to reflect commands and actions with target allocation and path planning, and quantifies the uncertainty by constructing a probabilistic ensemble model. Furthermore, the method has optimized the interaction mechanism and enhances the sample utilization based on the uncertainty quantification. We have also proposed an integration training method including pre–training and cross–training to enhance the training efficiency and stability. The experiment results have shown that our method achieves better episode returns, decision time, and confrontation win rate than the baselines, especially on large–size swarms. The influence of the adaptive frequency approach and ITM has been verified via ablation studies. Moreover, we have demonstrates the generalization of our method by applying a model trained under small–size swarms to the larger ones. After sufficient training, our method could emerge effective swarm confrontation strategies for agents.

### v2

@@ V-A Setting Up
Before analyzing the comparative results, we first introduce the detailed settings of simulations. Following the convention in [10, 25, 37], we generate the starting locations of pursuers and evaders in the rectangular area ($D_{1}=40m,D_{2}=20m,D_{3}=30m$). The target point of evaders is behind the pursuers. Three scenarios including ten pursuers versus ten evaders (termed as V10101010), fifteen pursuers versus fifteen evaders (termed as V15151515), and twenty pursuers versus twenty evaders (termed as V20202020) are considered in our experiment, all in the presence of four moving obstacles. In each scenario, we set up five different abilities of pursuers and evaders. The capture radius of pursuers $\rho_{c}$ is equalized from $0.6m$ to $1m$. The maximum velocity magnitudes of pursuers are all set to $0.5m/s$. The maximum velocity magnitudes of evaders are equalized from $0.6m/s$ to $1m/s$. The radius of the agents and obstacles are $0.2m$ and $0.5m$, respectively. Each obstacle keeps moving at a constant velocity over a certain time interval, while the magnitude $\lVert v_{o}\rVert\in[0.2,0.5]$ and direction $\psi_{o}\in[-\pi,\pi]$ of the velocity change randomly between every interval.
@@
We use the artificial potential field method with attractive and repulsive forces to find a motion vector for evaders. The preset target point exerts an attractive force in the direction of the vector between the target point and the evader. In addition, the pursuers, neighbors, and obstacles exert repulsive forces so that the evader can avoid being captured or collision. The repulsive forces decrease proportionally to the distance squared. The velocity is denoted as
$\begin{array}[]{c}\displaystyle v_{j}=\frac{p_{\text{tar}}-p_{j}}{\lVert p_{\text{tar}}-p_{j}\rVert}+\sum_{\begin{subarray}{c}w\in\left[U_{p},U_{e},O\right],\\ w\neq j\end{subarray}}{\frac{p_{j}-p_{w}}{\lVert p_{j}-p_{w}\rVert^{2}}},\end{array}$ (38)
where $p_{\text{tar}}$ is the position of the target point.
@@ V-C Comparison Analysis in Cross–Training
We further deploy the policy networks trained by the learning–based methods in $100$ different instances and compare them with the non–learning algorithms. We propose two additional evaluation metrics: decision time and confrontation win rate. Decision time refers to the total time taken by all pursuers for target allocation and path planning based on their observations. We take the average value after running different instances. The confrontation win rate refers to the ratio of the number of pursuers’ successes to the number of all instances. The average experiment results of $100$ instances in different sizes are shown in Table II. The table gathers the episode returns (Re.), decision time (Ti.), and confrontation win rate (W.R.) of all methods. We can observe that there is a relationship between the win rate and episode returns, which is calculated by the reward function in this study. Pursuers train their strategies to achieve higher returns, which leads to a greater win rate in the confrontation. Among these three non–learning algorithms, the heuristic algorithm has a longer decision time, although its returns and win rate are higher as the problem size increases. Due to the end–to–end approach, DRL has the shortest decision time, but its returns and win rate are much less than HRL. IMVE enhances the training efficiency of the method, but there is no additional improvement in its decision–making performance in the case of policy convergence. As a result, all three metrics for HRL and HRL/IMVE are similar in different scales. The results show that our method achieves better episode returns, decision time, and confrontation win rate than the baselines, especially on large–scale instances. The reason for this is that swarm confrontation in a dynamic obstacle environment has a high degree of uncertainty, and the baselines lack the ability to handle it. Our method optimizes the dynamic mechanism between the two layers based on the probabilistic ensemble model, which quantifies the uncertainty in the scenario.
@@ V-D Ablation Study for Adaptive Interaction Frequency
Moreover, we plot the value of $H$ for the adaptive method in different sizes and different episodes (0060606060) of cross–training in Fig. 7. The horizontal coordinate refers to the time steps in each episode, and the vertical coordinate refers to the value of $H$. In the early stage of training, the method decreases the interaction step $H$ to augment the training samples in the upper layer. Through constant training, the epistemic uncertainty in the lower layer decreases, so the interaction frequency becomes progressively lower. However, since aleatoric uncertainty cannot be eliminated, the increment in $H$ stabilizes after episode $15$. In addition, during time steps 50505050250250250250, the interaction frequency will be higher than the other time steps because the pursuers will meet with evaders and obstacles, and the aleatoric uncertainty will be higher at this time. Based on the adaptive frequency method, we can overcome the negative effects of uncertainty on cross–training, and finally achieve a favorable training effect.
@@ V-E Ablation Study on ITM
Figure 9: Generalization to larger–scale instances. (a) Four obstacles. (b) Eight obstacles.
@@ V-F Generalization to Larger–Size Swarms
We compare the model trained in V20 with the baselines in scenarios V25–V35 similar to Table II. The average experiment results are shown in Table III, which displays the episode returns (Re.), decision time (Ti.), and confrontation win rate (W.R.) testing in different problem sizes. As can be seen, HRL outperforms other baselines in all instances. To further investigate the generalization of our method, we utilize the trained model to solve instances with eight obstacles in Table III. An increase in the number of obstacles leads to a decrease in path rewards, but has a very small effect on confrontation win rate. The results demonstrate that the trained policies under small–size swarms are able to deal with the uncertainty in various–scale scenarios. Therefore, our method has a strong generalization on the swarm confrontation problem.
@@ V-G Emerged Coordinated Confrontation Behaviors
With sufficient training from scratch, the proposed HRL method can emerge effective target allocation and path planning strategies. Here, we visualize the results of trained networks with our method on one of the V10 instances in Fig. 10, and analyze the coordinated behaviors in each stage. Pursuers and evaders are labeled with "$X$" indicating their identification numbers. The figure illustrates the process of target allocation and path planning by pursuers through policy networks at different time steps. In the figure, the upper layer assigns targets based on pursuers’ and evaders’ attributes to maximize the total reward in (30). The lower layer further plans the safe paths to chase evaders while avoiding collisions with obstacles and neighbors.
@@
Time steps 50505050100100100100: After time step $50$, the pursuers begin to encounter evaders and obstacles, introducing more environmental uncertainty into decision–making. Since pursuers with a large radius have a wider capture range, the upper layer will prioritize mobilizing them to chase evaders with faster escape velocities, such as evader $4$ and evader $7$.
@@
Time steps 200200200200250250250250: After the evaders have escaped the initial round–up by the pursuers, the latter will choose to turn around and chase. The upper layer will assign the pursuers priority to evaders who are closer because they have a higher probability of being captured. Evaders may be hindered by the obstacles or neighbors ahead, resulting in a loss of velocities, and the pursuers eventually catch up with them.
Time steps 250250250250300300300300: Even though evader $9$ is approaching the target area, the others have been captured by pursuers. According to the rules in Section III-A, pursuers capture more than half of the evaders, which win the confrontation. Therefore, our trained models could deliver reasonably favorable solutions in the swarm confrontation.
Figure 11: Deployment in the real robot system.
V-H Deployment in Real–Robot System
To verify the adaptability of our method in the real world, we conduct experiments with an actual robot system shown in Fig. 11. We consider three pursuers versus three evaders in a static–obstacle scenario, where its setting is consistent with the definition in Section III-A. In the experiment, we maneuver agents by the ground control system with motion capture from Optitrack. By performing integration training in simulations, we can get the target allocation and path planning networks of pursuers, which can be deployed directly in real robots. Our algorithm receives messages from Optitrack, including the positions of all agents. Then the algorithm calculates the observation of each agent and outputs the position command with the above two networks, which is sent to the ground control system. Based on these networks, pursuers can capture all evaders while avoiding collision. The experiment’s success verifies that the designed method can accomplish the pursuit–evasion game through offline training and online decision–making.
@@ VI Conclusion
This paper has presented a guaranteed stable HRL method for hybrid decision spaces and high uncertainty in swarm confrontation. It has designed two–layer DRL networks to reflect commands and actions with target allocation and path planning, and quantifies the uncertainty by constructing a probabilistic ensemble model. Furthermore, the method has optimized the interaction mechanism and enhanced the sample utilization based on the uncertainty quantification. We have also proposed an integration training method including pre–training and cross–training to improve the training efficiency and stability. The experiment results have shown that our method achieves better episode returns, decision time, and confrontation win rate than the baselines, especially on large–size swarms. The influence of the adaptive frequency approach and ITM has been verified via ablation studies. Moreover, we have demonstrated the generalization of our method by applying a model trained under small–size swarms to the larger ones. After sufficient training, our method could emerge effective swarm confrontation strategies for agents and be deployed directly in the real–robot system.